OpenStudy (tester97):

Find an equation that models a hyperbolic lens with a = 16 inches and foci that are 40 inches apart. Assume that the center of the hyperbola is the origin and the transverse axis is vertical.

OpenStudy (tester97):

@ankit042 no one is helping ;(

OpenStudy (ankit042):

c=40 a=16 c^2 = a^2+b^2 so b =3 now as transverse axis is vertical y^2/16 - x^2/9 =1

OpenStudy (tester97):

i think that kinda makes sense

OpenStudy (ankit042):

k good!

OpenStudy (tester97):

thanks :)

OpenStudy (awkwardpanda):

do you remember the answer? @tester97